### v2

# Lanczos-Pascal approach to correlation functions in chaotic quantum systems

Abstract
We suggest a method to compute approximations to temporal correlation functions of few-body observables in chaotic many-body systems in the thermodynamic limit based on the respective Lanczos coefficients. Given the knowledge of these Lanczos coefficients, the method is very cheap. Usually accuracy increases with more Lanczos coefficients taken into account, however, we numerically find and analytically argue that the convergence is rather quick, if the Lanczos coefficients exhibit a smoothly increasing structure. For pertinent examples we compare with data from dynamical typicality computations for large but finite systems and find good agreement if few Lanczos coefficients are taken into account. From the method it is evident that in these cases the correlation functions are well described by a low number of damped oscillations.
Introduction. Understanding out-of-equilibrium quantum dynamics is one of the most fundamental and challenging questions in the field of statistical physics. It has received significant attentions, from both theoretical and experimental perspectives [1, 2, 3, 4]. Despite the substantial progress that has been made, many fundamental aspects of the problem remain unresolved, and it continues to be the focus of intensive research in the field.
Correlation functions play an essential role in characterizing out-of-equilibrium dynamics and bridge from theory to actual measurable physical phenomena. Over the past decades, the (re)introduction of the Eigenstate Thermalization Hypothesis (ETH) [5, 6, 7, 8] has provided a framework for studying these functions, particularly for the so-called autocorrelation functions. By postulating a particular structure of matrix elements of the observable in the eigenbasis of the Hamiltonian, the ETH explains the essential thermalization of autocorrelation functions after sufficiently long times. However, characterizing their full time evolution lies beyond the scope of the ETH.
To characterize the evolution of correlation functions, on the analytical side, except for trivial cases such as free fermions, general analytical predictions remain a major difficulty. In systems with conserved charge, hydrodynamics has been proven to be a useful tool to predict the long-time behavior of autocorrelation function [9, 10, 11]. However, it remains an assumption, and the extent of its applicability has yet to be fully established.
On the numerical side, the question is also involved due to the exponential complexity (in system size) of quantum systems. To tackle this question, various sophisticated approaches have been introduced. These range from perturbation theory [12, 13, 14], over quantum typicality [15, 16, 17], time-dependent density-matrix renormalization group methods [11, 18, 19] to quantum Monte-Carlo techniques [20, 21]; each with characteristic advantages and limitations, as well as distinct scopes in terms of application. However, when applied to large system sizes, most of them become quite involved and computationally expensive.
Another approach, the recursion method [22, 23, 24], significantly regained attention thanks to the recent Operator Growth Hypothesis (OGH) [25]. The latter constitutes a statement on the so-called Lanczos coefficients of an observable; quantities that fully characterize its dynamics in the thermodynamic limit. Eventually code bases such as [26] position this method along other well-established numerical techniques. Based on the OGH, the recursion method has been applied in various lines fo research [27]. Particularly it has been applied successfully in the context of correlation functions amounting to studies of dynamical quantities such as equilibration times [28, 29] and diffusion constants [30, 31]. The success in these applications triggered the question to address the full time evolution of correlation functions in this framework. Recently, in Ref. [32], an approach has been suggested to obtain correlation functions via an expansion in terms of pseudomodes, or Ruelle-Pollicot resonances [33, 34]. This method however presupposes the exact form of the Lanczos coefficients beyond the numerical scope, in performing a linear continuation of the latter motivated by the OGH.
In our work we put forward an approach to compute correlation functions on the basis of only the numerically accessible Lanczos coefficients that is both remarkably accurate yet very cheap. We provide analytical arguments that suggest a generic applicability of our method in systems in accord with the OGH. We augment these arguments by numerical investigations of pertinent toy models and specific quantum many-body spin chains, and benchmark using state-of-the-art numerical methods.
Setting. In this work, we focus on dynamics generated in operator space subject to the recursion method [22, 23, 24]. Concretely, consider an observable $|\mathcal{A})$ regarded as a vector in operator space. The time evolution is implemented by the Liouvillian superoperator $\mathcal{L}=\left[\mathcal{H},-\right]$, where $\mathcal{H}$ denotes the systems’ Hamiltonian operator. The Lanczos algorithm orthonormalizes the set of vectors $\{\mathcal{L}^{j}|\mathcal{A})\}_{j=0}^{\infty}$ in a Gram-Schmidt procedure. Concretely, the algorithm is initialized by a normalized seed vector $|\mathcal{O}_{0})=|\mathcal{A})/\sqrt{(\mathcal{A}|\mathcal{A})}$, where $(\mathcal{A}_{1}|\mathcal{A}_{2})=\text{tr}\{\mathcal{A}_{1}^{\dagger}\mathcal{A}_{2}\}/\text{tr}\{\mathbb{1}\}$ denotes the Hilbert-Schmidt scalar product. For $n>0$ it then iteratively computes $|\tilde{\mathcal{O}}_{n})=\mathcal{L}|\mathcal{O}_{n-1})-b_{n-1}|\mathcal{O}_{n-2})$ and normalizes $|\mathcal{O}_{n})=|\tilde{\mathcal{O}}_{n})/(\tilde{\mathcal{O}}_{n}|\tilde{\mathcal{O}}_{n})^{1/2}=:|\tilde{\mathcal{O}}_{n})/b_{n}$, where by convention $b_{0}=0$. The quantities $b_{n}$ are the so-called Lanczos coefficients and the Krylov vectors $|\mathcal{O}_{n})$ span the Krylov space in which the Liouvillian takes tridiagonal form. In order to address dynamical quantities of a system at hand, we define auxiliary functions
$\varphi_{n}(t):=i^{-n}(\mathcal{O}_{n}|\mathcal{O}(t)),$ (1)
which inherit from the Liouvillian the discrete Schrödinger equation
$\partial_{t}\varphi_{n}(t)=b_{n}\varphi_{n-1}(t)-b_{n+1}\varphi_{n+1}(t).$ (2)
For later reference we note that Eq. (2) may be cast into a matrix equation
$\dot{\vec{\boldsymbol{\varphi}}}=\textbf{L}\cdot\vec{\boldsymbol{\varphi}},\quad\text{where}\ \textbf{L}_{nm}=b_{n}\delta_{n,m-1}-b_{n-1}\delta_{n,m+1}.$ (3)
Imposing the initial condition $\varphi_{n}(0)=\delta_{n,0}$, solving Eq. (3) amounts to the computation of the infinite-temperature autocorrelation function of the observable of interest, as $\mathcal{C}(t)=(\mathcal{O}_{0}(t)|\mathcal{O}_{0})=\text{tr}\{\mathcal{O}(t)\mathcal{O}\}/\text{tr}\{\mathbb{1}\}=\varphi_{0}(t)$. Conversely, other entries of $\vec{\boldsymbol{\varphi}}$ represent correlation functions between the initial operator and other operators in Krylov space, cf. Eq. (1).
In principle, solving the matrix equation (3) allows for an exact computation of dynamical quantities such as the autocorrelation function, given an infinite amount of Lanczos coefficients. However, typically only a low, two-digit number of $b_{n}$ is numerically feasible, depending on the structure of system and observable alike. To cope with this limitation, the Lanczos coefficients are sometimes extrapolated in accord with the operator growth hypothesis (OGH) [25]. The latter states that, for few-body observables in chaotic quantum many-body systems that do not have overlap with any conserved quantity, the Lanczos coefficients asymptotically grow linear. Numerically, this behavior has been affirmed in several relevant settings, see e.g. [25, 35, 36, 30, 31]. It is worth noticing, however, that any truncation of L, albeit at a high dimension, will give rise to echoes, i.e. unphysical revivals of the correlation function, see e.g. [30].
[…]
$\mathcal{A}_{r}=\frac{1}{b_{1}}\prod_{n=1}^{r-1}\left(\frac{b_{n}}{b_{n+1}}\right)^{(-1)^{n}}\sqrt{\left(\frac{b_{r}}{b_{r+1}}\right)^{(-1)^{r}}}$ (4)
This estimate has been suggested in Ref. [37]. It is expected to yield better results with increasing $r$. Let $R$ denote the position after which Eq. (4) has converged suitably. Note that this involves $R+1$ Lanczos coefficients. Next, we infer $D$ as the smallest integer such that $\left(b_{1}/b_{R+1}\right)^{D}\leq\delta$, with some suitably chosen, small $\delta$ (below we employ $\delta=1/5$). This second condition requires the Lanczos coefficients to grow. With both $R,D$ and $R+D<n_{\max}$ we define
$s_{n}=\begin{cases}0&0\leq n\leq R-2\\ (-1)^{n+R+1}\binom{D}{n-R-1}&R-1\leq n\leq R+D-1.\end{cases}$ (5)
For reasons that shall be explained later, we refer to $D$ as the ”order” of the LP curve. If both above conditions, involving $R$ and $D$, respectively, are met within the accessible $b_{n}$, we expect the LP approach to yield rather accurate results. This conforms to all examples presented below. Note that the non-zero entries in Eq. (5) are the Pascal numbers with alternating sign, motivating the name of our scheme.
Eventually, we approximate the actual dynamics of $\vec{\boldsymbol{\varphi}}$, as imposed by L, by solving the simple matrix equation
$\dot{\vec{\boldsymbol{\psi}}}=\textbf{S}\cdot\vec{\boldsymbol{\psi}},$ (6)
where $\vec{\boldsymbol{\psi}}(t)=(\psi_{0}(t),\psi_{1}(t),\dots,\psi_{N-1}(t))^{T}$ with $N:=R+D$ and
$\textbf{S}=\begin{pmatrix}0&-b_{1}&0&0&\dots\\ b_{1}&0&-b_{2}&0&\dots\\ 0&b_{2}&0&\ddots&\dots\\ \vdots&\vdots&\ddots&\ddots&-b_{N-1}\\ b_{N}s_{0}&\dots&b_{N}s_{N-3}&b_{N}s_{N-2}+b_{N-1}&b_{N}s_{N-1}\end{pmatrix},$ (7)
and the initial condition $\psi_{n}(0)=\delta_{n0}$. Similar approaches to truncations of the Liouvillian have been introduced in [38, 39]. The LP approach now asserts that $\psi_{n}(t)\approx\varphi_{n}(t)$ for $n=0,\dots,N-1$. Consequently, in analogy to Eq. (1), solutions of the simple $N\times N$ matrix equation (6) give rise to approximations of several relevant dynamical quantities ranging from autocorrelations to expectation values in quench settings. Concretely, we have
$\psi_{0}(t)$ LP${\approx}\left(\mathcal{O}_{0}(t)|\mathcal{O}_{0}\right)\propto\text{Tr}\{\mathcal{O}_{0}(t)\mathcal{O}_{0}\},$ (8)
$\psi_{1}(t)$ LP${\approx}\left(\mathcal{O}_{1}(t)|\mathcal{O}_{0}\right)\propto\text{Tr}\{\left[\mathcal{H},\mathcal{O}_{0}\right](t)\mathcal{O}_{0}\}\dots$ (9)
It is worth noting that by construction, solutions $\psi_{k}(t)$ are a superposition of at most $N$ damped exponentials. In this sense, LP curves are simple for low $N$.
Numerics. Before turning to physical models, we consider settings in which the connection between the Lanczos coefficients and the analytical form of the autocorrelation function is known. To this end, we study the toy model given by $b_{n}=1+\alpha(n-1)$ for $0\leq\alpha\leq 1$. In the cases where $\alpha=0$ and $\alpha=1$, the autocorrelation functions are known to be $C(t)=J_{1}(2t)/t$ and $C(t)=\text{sech}(t)$ respectively [37]. Here $J_{n}(x)$ denotes the Bessel functions of first kind. The long-time behavior of the former is known to follow a power-law, i.e. an approximation with few exponentials is inherently flawed. We note that in line with our construction, for constant $b_{n}$, no $D$, as described below Eq. (4) can be inferred. Here we focus on the two cases $\alpha=0.1,1$ specifically. While for $\alpha=1$ we may compare with the analytical form of the autocorrelation function, for $\alpha=0.1$ we benchmark the LP scheme against the solution of the matrix equation (3) with dimension 2000  [^1] As explained in the introduction, solutions of Eq. (3) inevitably features echoes. Here we therefore only consider times where the correlation function has decayed suitably but well before the echo sets in..
Throughout the paper, we determine $R$ as the position after which $\mathcal{A}_{r}$ varies less that $1/200\max_{j}\mathcal{A}_{j}$ from subsequent areas. Further, we set $\delta=0.2$. In Fig. 1 we show the LP scheme at the example of the toy model. Here, the area $\mathcal{A}_{R}$ has converged in the above sense at $R=1$ for $\alpha=0.1$ and at $R=9$ for $\alpha=1$. The quantities $D$ are determined to be 19 and 1, respectively. We find that in both cases the LP method gives rise to approximations to the autocorrelation functions with strong agreement to the actual dynamics. For the toy model with $\alpha=0.1$ we also depict the respective LP curve with $(R,D)=(19,1)$ in order for a fair comparison such that in both scenarios the curves consist of the same number of damped oscillations. Unlike for $(R,D)=(1,19)$, the curve eventually fails to capture the true dynamics and deviates visibly indicating the importance of a suitable order in the LP scheme.
Figure 1: LP method for a toy model. Area estimator $\mathcal{A}_{r}$ and the autocorrelation function obtained by the LP method and benchmark (inset: log-plot) for $\alpha=0.1$ (left) and $\alpha=1$ (right).
Figure 2: LP scheme for the autocorrelation function. For each system and each observable we depict the respective first $n_{\max}$ Lanczos coefficients $b_{n}$, the area estimate $\mathcal{A}_{r}$ along with the point of convergence $R$ and both the autocorrelation function obtained by the LP method and DQT. From left to right: Fast-mode observable in the chaotic Ising chain with $B_{z}=0.5$, $[(a),(b),(c)]$, slow-mode observable in the chaotic Ising chain with $B_{z}=0.5$, $[(d),(e),(f)]$, energy current in the two-dimensional Ising chain, $[(g),(h),(i)]$, local $x$ component in the spin-bath model, $[(j),(k),(l)]$, fast-mode observable in the tilted-field Ising model with $B_{z}=2.0$, $[(m),(n),(o)]$.
As a first quantum system, we consider as an example of a generic chaotic quantum system a tilted-field Ising model (TFIM) with the Hamiltonian $\mathcal{H}=\sum_{k}h_{k}$, where
$h_{k}=\sigma_{k}^{z}\sigma_{k+1}^{z}+\frac{B_{x}}{2}\left(\sigma_{k}^{x}+\sigma_{k+1}^{x}\right)+\frac{B_{z}}{2}\left(\sigma_{k}^{z}+\sigma_{k+1}^{z}\right),$ (10)
with $B_{x}=1.05$ and two different field strengths $B_{z}=0.5,2.0$. In this model, we investigate several different observables, altogether advocating the generality of our findings. In the main part of this paper, we consider density waves of the energy with both short and long wavelengths,
$\mathcal{O}_{q}\propto\sum_{k}\cos(qk)h_{k}\quad\text{for}\ q=\pi,\pi/14.$ (11)
Here, we consider up to $n_{\text{max}}=40$ Lanczos coefficients. See App. C for further examples in the tilted-field Ising model. Next, we include into our study a two-dimensional transverse-field Ising model,
$\mathcal{H}=\sum_{\langle ij\rangle}\sigma_{i}^{x}\sigma_{j}^{x}+\sum_{i}\sigma_{i}^{z},$ (12)
where by the notation $\langle ij\rangle$ the sum is restricted to adjacent sites on the two-dimensional lattice. As an observable of interest we consider the energy current
$\mathcal{J}_{E}^{2d}=\sum_{\langle ij\rangle|i<j}\sigma_{i}^{x}\sigma_{j}^{y}-\sigma_{j}^{x}\sigma_{i}^{y}.$ (13)
This setting has recently been studied in a related context, from where also the first 21 Lanczos coefficients are taken [32].
Figure 3: LP scheme of higher correlators. Correlations functions between the first two Krylov states of the respective observable obtained by the LP method and DQT. From left to right: Fast mode and slow mode of the energy density-wave operator in the tilted-field Ising model with $B_{z}=0.5$, energy current in the two-dimensional Ising model, $x$-component of the local spin in the spin-bath model.
Lastly, we study an open-system setup, where a single spin is coupled to a large Ising bath. Concretely, we study
$\mathcal{H}=\lambda\sigma_{0}^{z}\sigma_{1}^{z}+\sum_{j}\sigma_{j}^{z}\sigma_{j+1}^{z}+B_{x}\sigma^{x}_{j}+B_{z}\sigma_{j}^{z},$ (14)
with $(\lambda,B_{x},B_{z})=(2,1,0.5)$. Here, the observables of interest is the respective $x$ component of the single spin, i.e. $\sigma_{0}^{x}$, for which we determine $n_{\max}=40$ Lanczos coefficients.
In order to benchmark the dynamics we obtain from the LP method, we compare with state-of-the-art numerical techniques on the basis of dynamical quantum typicality (DQT)[41, 17]. For the tilted-field Ising chain, we consider a finite system with length $L=28$, for the transverse Ising model we study a $5\times 5$ lattice and for the Ising bath in (14) we have $L=28$. For all systems we employ periodic boundary conditions. The Lanczos coefficients $b_{n}$ and thus the dynamics generated by the LP approach always concern the thermodynamic limit, i.e. $L\rightarrow\infty$.
In Fig. 2 we depict the autocorrelation functions for the models and observables introduced above. We find that whenever the LP approach is feasible, i.e. the area $\mathcal{A}_{R}$ has converged in the above sense and we may infer $D$, the LP curve is practically indistinguishable from the actual data. It is worth mentioning that this agreement is present, regardless of the feature of the autocorrelation functions, i.e. few or many oscillations or mere decay. Deviations set in on a scale that is negligible for most practical purposes. For the tilted-field Ising model with a strong perpendicular field $B_{z}$, here $B_{z}=2.0$, the areas $\mathcal{A}_{r}$ do not converge within our numerical scope, hence we cannot infer an LP curve subject to the construction introduced above and our $n_{\max}$. It is worth noticing that, for a magnetic field of this strength, the finite system leaves the chaotic regime, see e.g. [29]. To highlight the inapplicablity of the LP method in this scenario, we depict in Fig. 2 $m)$ two curves, each consisting of the maximal number of damped oscillations allowed by the LP scheme. Concretely, we show cases where $(R,D)=(40,1)$ and $(31,10)$ respectively.
Further, via the LP approach we may address other correlation functions. For the previous examples with convergent areas, we show in Fig. 3 the LP approximations $\psi_{1}(t)$ to higher correlators. As with the autocorrelator and $\psi_{0}(t)$ respectively, the agreement of the LP method is remarkable with deviations only visible on the log-scale.
For all observables that feature smoothly increasing Lanczos coefficients $b_{n}$, see Figs. 1 and 2 $[(c),(f),(i),(l)]$, we find the area estimator $\mathcal{A}_{r}$ to converge within our numerical scope and eventually yielding remarkably accurate approximations via the LP method. Conversely, for cases with ragged $b_{n}$ such as for the fast-mode observable in the TFIM with strong field, see Fig. 2 $(o)$, no LP curve may be obtained for our $n_{\max}$. Applicability of the Lanczos-Pascal method. The applicability of the LP scheme is closely tied to the structure of the Lanczos coefficients $b_{n}$ themselves, see also App. A,B. In order for the LP scheme to apply, the respective areas need to converge. Considering
$\frac{\mathcal{A}_{r+1}}{\mathcal{A}_{r}}=\sqrt{\left(\frac{b_{r}b_{r+2}}{b_{r+1}^{2}}\right)^{(-1)^{r}}}$ (15)
we infer that for convergent areas $\mathcal{A}_{r}$, the Lanczos coefficients need to become sufficiently smooth, i.e. $b_{r}b_{r+2}/b_{r+1}^{2}\approx 1$. It is worth stressing that for systems within the scope of the operator growth hypothesis, the latter condition may be expected to be generically fulfilled. Revisiting the Lanczos coefficients for the settings considered in this paper in Fig. 2 and Fig. 3, precisely the cases with smoothly increasing $b_{n}$ each give rise to accurate descriptions by means of the LP method. In essence, the comparatively weak requirement posed in Eq. (15) strongly suggests the generic applicability of the LP method in chaotic quantum many-body systems.
Conclusion. In this paper, we introduced the Lanczos-Pascal approach as a scheme based on the recursion method to approximate dynamics in quantum many-body systems. Taking into account only few, numerically accessible Lanczos coefficients, the method is very cheap. We formulated criteria based on the Lanczos coefficients of an observable in order to a priori assess whether the proposed method is applicable in a given setting. We numerically checked the method for a multitude of systems, ranging from a toy model, over pertinent chaotic Ising chains in both one and two dimension to an open system spin-bath setup. In all cases, the LP approximations agree remarkably well with state-of-the-art benchmarks obtained by means of dynamical quantum typicality, given that predefined conditions on the Lanczos coefficients indicated the LP to be feasible.
We augment our numerical findings by analytical considerations that link the convergence of the LP scheme to structures in the respective Lanczos coefficients. Crucially, in the context of few-body observables in chaotic quantum many-body systems, these considerations are supplemented by the operator growth hypothesis (OGH). With the latter putting forward an eventual linear growth of the Lanczos coefficients, our arguments suggest the generic applicability of our scheme. In App. B we provide similar arguments for the case of finite frequencies $\omega\neq 0.$
Moreover, the feasibility of the LP approach has further conceptual implications. By construction the LP approach involving $m$ Lanczos coefficients will yield approximations that each consist of at most $m$ damped exponentials. In this interpretation, curves that yield LP curves with low $m$ may be regarded as simple. This addresses nature’s apparent soft spot for simple dynamics, oftentimes appearing to favor certain classes of relaxation over others.
Note added. While this manuscript was in preparation, a related study appeared [39].
Acknowledgement. We thank Christian Bartsch, Robin Heveling and Mats Lamann for fruitful and valuable discussions. This work has been funded by the Deutsche Forschungsgemeinschaft (DFG), under Grant No. 531128043, as well as under Grant No. 397107022, No. 397067869, and No. 397082825 within the DFG Research Unit FOR 2692, under Grant No. 355031190.
Appendix A The Lanczos-Pascal scheme as a derivative in Krylov space
The applicability of the Lanczos-Pascal scheme is closely related to the structure of the wave vector $\vec{\boldsymbol{\varphi}}$. Assuming that $\vec{\boldsymbol{\psi}}$ captures the first $N$ of $\vec{\boldsymbol{\varphi}}$ precisely, i.e. $\varphi_{n}(t)=\psi_{n}(t)$ for $n=0,...,N-1$, combining Eq. (3) and Eq. (6) we may write
00 0 $=\left(\textbf{S}-\textbf{L}\right)\vec{\boldsymbol{\varphi}}\ \Longleftrightarrow\ \vec{s}\cdot\vec{\boldsymbol{\varphi}}=0.$ (16)
Consequently, the existence of a matrix S that generates the first $N$ entries of $\vec{\boldsymbol{\varphi}}$ correctly is equivalent to a vector $\vec{s}$, orthogonal to the wave function. Further, with the vector $\vec{s}$ as defined in Eq. (5) the scalar product represents a discrete spatial derivative of the wave function. Hence whenever the wave vector $\vec{\boldsymbol{\varphi}}$ with suitably smooth with respect to its components we expect the LP scheme to yield appropriate results. In this interpretation the quantity $D$ corresponds to the order of the derivative. See Fig. 4 for an illustration of the wave vector $\vec{\boldsymbol{\varphi}}$ related to the fast-mode observable in the tilted-field Ising chain with $B_{z}=0.5$ for an impression of a smooth wave vector.
Figure 4: Wave vector $\vec{\boldsymbol{\varphi}}$ related to the fast-mode of energy-density wave in the tilted-field Ising model with $B_{z}=0.5$ for various instances of time. Differently to the analysis in the main text, for simplicity we resort to solving the matrix equation (3).
While this consideration yields an intuitive picture of the inner workings of the LP method, it clearly does not provide solid justification for the applicability of the latter. This is so because the smoothness of $\vec{\boldsymbol{\varphi}}$ is phenomenologically demonstrated in Fig. 4, but not generally mathematically established based on properties of $b_{n}$. The latter will be done below.
Appendix B Smoothness of the wave vector
Consider the Laplace transform of Eq. (3) to find
$\begin{pmatrix}-\varphi_{0}(0)\\ -\varphi_{1}(0)\\ \vdots\end{pmatrix}=\begin{pmatrix}-1\\ 0\\ \vdots\end{pmatrix}=\textbf{L}\begin{pmatrix}\widetilde{\varphi}_{0}(0)\\ \widetilde{\varphi}_{1}(0)\\ \vdots\end{pmatrix},$ (17)
where by $\widetilde{\varphi}_{k}(0)$, subsequently abbreviated by $\widetilde{\varphi}_{k}$, we denote the Laplace transform of the $k$th entry of the approximated wave vector $\vec{\boldsymbol{\varphi}}$ evaluated at zero. Note that $\widetilde{\varphi}_{0}$ equals the area under the curve, i.e, the quantity that we aim to estimate by $\mathcal{A}_{r}$ in Eq. (4). Equation (17) may be solved ”from top to bottom”. Doing so, every $\widetilde{\varphi}_{n}$ with $n$ even is determined up to a multiplicative factor $\widetilde{\varphi}_{0}$. Concretely,
$\widetilde{\varphi}_{2n}/\widetilde{\varphi}_{0}$ $=1,\frac{b_{1}}{b_{2}},\frac{b_{1}b_{3}}{b_{2}b_{4}},\dots$ (18)
$\widetilde{\varphi}_{2n+1}$ $=\frac{1}{b_{1}},\frac{b_{2}}{b_{1}b_{3}},\frac{b_{2}b_{4}}{b_{1}b_{3}b_{5}},\dots,$ (19)
Eventually, we may write
$\mathcal{A}_{r}$ $=\widetilde{\varphi}_{0}\sqrt{\left(\frac{\widetilde{\varphi}_{r-1}\widetilde{\varphi}_{r+1}}{\widetilde{\varphi}_{r}^{2}}\right)^{(-1)^{r}}}.$ (20)
Consequently, convergence of $\mathcal{A}_{r}$ implies $\mathcal{A}_{r}\approx\widetilde{\varphi}_{0}$. Thus, for $r$ in the regime where $\mathcal{A}_{r}$ has converged, say at $r\geq R$, the square root term in Eq. (20) has to be very close to 1, independent of the specific $r$. This in turn implies that the sequence of the components $\widetilde{\varphi}_{r}$ has become smooth. This eventually amounts to the Laplace transformation of Eq. (16) at zero, i.e.
$\vec{s}\cdot\vec{\widetilde{\boldsymbol{\varphi}}}=0,$ (21)
typically being approximately fulfilled for pertinent $\vec{s}$, as this construction amounts to taking a discrete derivative of order $D$ of a smooth sequence. Further, we extend our arguments to nonzero arguments of the Laplace transformed quantities. Since the $\tilde{\varphi}_{n}(s)$ do not feature any poles on the right half complex plain, we may restrict our analysis to the imaginary axis, i.e. $\omega\neq 0$. (The Laplace transform thus amounts to a Fourier transform). On the latter, the Laplace/Fourier transform of (3) becomes
$-\vec{\boldsymbol{\varphi}}(0)=\left(-i\omega\ \mathbb{1}+\textbf{L}\right)\vec{\widetilde{\boldsymbol{\varphi}}}(\omega)$ (22)
or similarly
$i\omega\widetilde{\varphi}_{n}(\omega)=b_{n+1}\widetilde{\varphi}_{n+1}(\omega)-b_{n}\widetilde{\varphi}_{n-1}(\omega).$ (23)
Upon defining
$\widetilde{\varphi}_{n}(\omega)=:\underbrace{f_{n}(\omega)}_{f_{n}}\underbrace{\widetilde{\varphi}_{n}(0)}_{\widetilde{\varphi}_{n}},$ (24)
we may link Eq. (23) to the setting where $\omega=0$ and write
$i\omega f_{n}\widetilde{\varphi}_{n}$ $=(f_{n+1}-f_{n-1})b_{n+1}\widetilde{\varphi}_{n+1}$ (25)
$\Leftrightarrow\left(f_{n+1}-f_{n-1}\right)$ $=i\omega f_{n}\left(\frac{\widetilde{\varphi}_{n+1}}{\widetilde{\varphi}_{n}}b_{n+1}\right)^{-1}.$ (26)
As argued around Eq. (20), for $n\geq R$, we have $\widetilde{\varphi}_{n+1}/\widetilde{\varphi}_{n}\approx 1$. As argued around Eq. (15), the $b_{n}$ form a smooth sequence in the same regime. With this, we take the continuum limit of Eq. (26) to find
$2f^{\prime}(n)$ $\approx i\omega f(n)/b(n)$ (27)
$\Rightarrow f(n)$ $=\exp\left(\frac{i}{2}\omega\int_{n_{0}}^{n^{\prime}}\frac{1}{b(n^{\prime})}\text{d}n^{\prime}\right).$ (28)
Figure 5: LP method in the tilted field Ising model (TFIM). The two-local observable $\mathcal{B}$ (left) and energy current $\mathcal{J}_{E}$ (right) in the tilted-field Ising model, given by Eq. (29) and (30) respectively. From top to bottom: Autocorrelation function obtained by the LP method and DQT (inset: log-plot), the correlation between the first two Krylov states of the respective observable, again by means of the LP method and DQT (inset: log-plot), the area estimate $\mathcal{A}_{r}$ with convergence position $R$ and the respective Lanczos coefficients $b_{n}$.
Considering finite frequencies $\omega\neq 0$, we infer from Eq. (28) that for Lanczos coefficients $b_{n}$ that grow according to the OGH, for small frequencies $\omega$, the functions $f_{n}$ oscillate slowly. Hence this allows to connect the picture with the $\omega=0$ setting and therefore implies that the LP scheme is also applicable in this regime. Conversely, for Lanczos coefficients that do not grow, the area $\mathcal{A}_{r}$ is constant for all $r$. However, it is evident from Eq. (28) that already at small frequencies the functions $f_{n}$ will oscillate strongly and therefore the scenario may not simply be mapped to the $\omega=0$ case. This behavior is perfectly in line with the LP technique explained in the main part of the paper. In the case of constant Lanczos coefficients the autocorrelation function is known to be a Bessel function, see [37], and further there is no such $D$ fulfilling $(b_{1}/b_{R+1})^{D}<1$ for any $R$.
Appendix C Further examples in the tilted-field Ising model
In the tilted-field Ising model we include further to our study the two-local observable $\mathcal{B}$ and the energy current $\mathcal{J}_{E}$. Concretely these operators are given by
$\mathcal{B}$ $\propto\sum_{k}J\sigma_{k}^{x}\sigma_{k+1}^{x}-\sigma_{k}^{z},$ (29)
$\mathcal{J}_{E}$ $=B_{x}\sum_{k}\sigma_{k}^{y}\left(\sigma_{k+1}^{z}-\sigma_{k-1}^{z}\right),$ (30)
with $J=1.05$, for which we compute $n_{\max}=40$ and $n_{\max}=44$ Lanczos coefficients respectively. In Fig. 5 we depict the LP curves for both observables for both $\psi_{0}(t)$ and $\psi_{1}(t)$ as well as the area estimator $\mathcal{A}_{R}$. For the two-local observable $\mathcal{B}$ we find a suitably converged area at $R=18$ and a necessary order $D=2$. For the energy current we have $(R,D)=(35,1)$. Again, the LP scheme gives rise to approximation that are distinguishable from the DQT data to the naked eye. Here, the functions $\psi_{0}(t)$ and $\psi_{1}(t)$ both only consist of at most 20 and 36 exponentially damped constituents for the two observables respectively.